OpenStudy (anonymous):

Help Please! ? :) Just Need The Steps :) 3 + 1 = 6 ------ -------- -- m-4 3(m-4) m

OpenStudy (kewlgeek555):

Well, I think that this problem is stating that m=6. If m equals 6 then you can solve for the other ones. 6 - 4 = 2. 3(6 - 4) = 6. So \[\frac{ 2 }{ 6 }\] I think. I am just in eighth grade. At least I tried to help.

OpenStudy (anonymous):

The Answer iS 9, iJust Need To Know How They Got Tht

OpenStudy (dumbcow):

first combine fractions on left \[\frac{3}{m-4} + \frac{1}{3(m-4)} = \frac{9}{3(m-4)}+\frac{1}{3(m-4)} = \frac{10}{3(m-4)}\] then cross-multiply \[\frac{10}{3(m-4)} = \frac{6}{m}\] \[10m = 6*3(m-4)\] \[10m = 18m - 72\] \[8m = 72\] \[m=9\]

OpenStudy (anonymous):

Thank You So Much :)

OpenStudy (dumbcow):

yw

OpenStudy (kewlgeek555):

Now I feel dumb. Well, I bet this is like Algebra III or somethin'. I'm just in Math III. Sigh.

OpenStudy (anonymous):

Lol... Its Algebra 1

OpenStudy (kewlgeek555):

Oh! I am going there next year! :3

OpenStudy (anonymous):

Nice :)